OpenStudy (liv33):

Solve: 3/5 + d is less than or equal to 13/5

OpenStudy (samigupta8):

Can you make an equation for this?

OpenStudy (mariamikayla):

3/5+d <= 13/5 we bring to the same denominator every therm of the equation 3+5d <=13 5d<=13-3 5d <=10 d<=10/5 d<=2 => d e (-infinity;2]

OpenStudy (liv33):

Thanks!